# Copyright in AI Pre-Training Data Filtering: Regulatory Landscape and Mitigation Strategies

### v1

@@ 2.4.1 Japan
Japan’s revised Copyright Act (Article 30-4, effective January 1, 2019) allows broad rights to ingest and use copyrighted works for any type of information analysis, including AI training, for both commercial and non-commercial purposes [privacyworld2024japan]. Unlike the EU, Japan allows commercial use and reportedly even permits training on ”content obtained from illegal sites,” with Japan’s Education Minister indicating AI companies can use ”whatever they want” for training [privacyworld2024japan].
@@ 2.4.2 Singapore
Singapore’s Copyright Act Amendment of 2021 introduced Section 244, allowing use of copyrighted works for ”computational data analysis” including commercial use, provided users have ”lawful access” to the content [techpolicy2024asia].
@@ 2.4.3 China
China’s ”Interim Measures for the Management of Generative AI Services” (July 2023) requires that generative AI services ”respect intellectual property rights,” with Article 7.2 stipulating that AI providers must not infringe others’ IP rights [lexology2024apac, techpolicy2024asia]. China also requires transparency about training data use and has shown willingness to impose secondary liability on AI platforms for copyright infringement [iic2025copyright, globallegal2025china].
@@ 2.4.4 South Korea
Korea requires permission for using copyrighted works for AI training, as emphasized in Copyright Commission guidelines published in December 2023. Courts have ruled that web scraping infringes database rights under copyright law [lexology2024apac, jtip2021copyright].
@@ 3.1 Pre-Training License Collection and Content Filtering
Modern large language models require training datasets containing billions of text tokens, images, and other media scraped from publicly accessible web sources. The sheer volume of this data collection makes traditional copyright clearance mechanisms practically impossible to implement.
@@
Regulatory Gap: the US Copyright Office noted, ”it is not practically possible to obtain licenses for the volume and diversity of content necessary to power cutting-edge systems” [uscopyrightoffice2025]. New websites, blog posts, images, and other copyrighted materials appear online continuously, often before rights holders can implement or update restriction mechanisms. This temporal lag between content publication and rights assertion creates unavoidable exposure windows where AI companies may unknowingly incorporate restricted content into their training datasets.
@@ 3.2 Verification Mechanisms Gaps
Regulatory Gap: On the other side, the regulatory verification mechanism still has its gaps. For example, the EU AI Code of Practice exemplifies this verification deficit. While Measure 1.4 [eu2024codeofpractice, Copyright Chapter, Measure 1.4] requires GPAI providers to describe their data filtering methodology in Technical Model Documentation, the Code provides no technical implementation guidelines for verification systems. Companies must document their processes, but no standardized audit mechanism exists to verify that these documented processes were actually followed or that they effectively prevent copyright infringement. Unless the complaint has been filed and then the process of verification can take place. This highlights the reactive approach to pre-training data filtering rather than preventive.
@@
An example of mentioned above challenges is The New York Times’ lawsuit against Microsoft and OpenAI [nyt2023openai] which illustrates the most serious category of potential copyright infringement. The Times alleges that AI models were trained on wholesale copying of Times articles, and that the resulting AI capabilities directly compete with the newspaper’s core business model. The Times demonstrated instances where AI models could reproduce content substantially similar to specific Times articles, suggesting that training data patterns were being memorized rather than truly transformed. This raises significant problems as the training data used can’t be tracked as well as after training it is impossible to access the dataset.
These three challenges correspond to the open research questions identified in Stanford’s TAIG Database [stanford2024taig] discussed before, demonstrating that the enforcement gaps we have identified reflect broader, unsolved problems in technical AI governance.
@@ 4 Existing Mitigation Strategies
One more solution is the ”pay per crawl” model [cloudflare2024paypercrawl], which allows content creators to have control over who accesses their work. If a creator wants to block all AI crawlers from their content, they should be able to do so. If a creator wants to allow some or all AI crawlers full access to their content for free, they should be able to do that, too. Moreover, pay per crawl grants domain owners full control over their monetization strategy. They can define a flat, per-request price across their entire site, meaning they get paid if an AI company wants to scrape the data from their website. Companies such as CloudFlare [cloudflare2024paypercrawl] have begun implementing and developing this approach, offering platforms for content creators to manage crawler access and monetize their data.
@@
Another solution is Named Entity Recognition (hereinafter NER) [li2022survey] which is used in AI Chatbots and security tools. It is a natural language processing technique that identifies and classifies specific entities in text into predefined categories such as: person names, organizations, locations, chemical compounds, works of art. NER could be adapted to identify copyrighted content by: detecting author names, publication titles, or publisher names; identifying trademarked product names or brand references; recognizing specific works (book titles, article headlines, song names); matching against databases of known copyrighted entities. However, NER alone is insufficient for copyright filtering because copyright protects expression, not just entity names. For example, a text could discuss ”The New York Times” without containing any Times content. Conversely, copyrighted text might not contain obvious entity markers, and, moreover, NER can’t detect paraphrased or derivative content.
@@
Also perceptual hashing [sun2022deep, tang2024survey] is a great method that is already in use by many companies. Unlike cryptographic hashes (which change completely with any tiny modification), perceptual hashes create a ”fingerprint” that remains similar even when content is slightly altered (resized, compressed, cropped, color-adjusted). The system can then compare these hashes to identify matching or near-matching content. While this technology is widely deployed by platforms like YouTube and Facebook [titlow2016youtube] to identify copyrighted material after upload, it remains underutilized in the AI training data context. AI companies could theoretically check billions of scraped images and videos against copyright databases using perceptual hashing before including them in training datasets, but there’s little evidence of systematic implementation, likely due to computational costs, incomplete database coverage, and legal ambiguity about whether using copyrighted images for training (without reproducing them in outputs) constitutes infringement.
@@
Have I Been Trained [HaveIBeenTrained] is a searchable database that allows artists and creators to check whether their images appear in major publicly documented AI training datasets like LAION-5B and LAION-400M, which were used to train models such as Stable Diffusion. Users can search by uploading images, entering artist names, or providing URLs to discover if their work has been included in training data, and can request opt-out from future dataset versions. However, the tool has significant limitations: it only covers datasets that companies have publicly disclosed (excluding proprietary datasets like OpenAI’s), works retrospectively rather than preventing initial scraping, provides no legal enforcement mechanism, covers only images rather than text or code, and relies entirely on voluntary compliance from AI companies. This shows the reactive nature of current copyright protection mechanisms, creators can discover violations only after their work has already been scraped and used for training, and opt-out requests don’t affect models already trained on the data, highlighting the gap between post-collection transparency tools and the proactive pre-training filtering mechanisms needed to prevent copyright infringement before it occurs.
@@
InnerProbe [ma2024innerprobe] presents a post-training framework for identifying copyright-related influences in Large Language Model outputs by leveraging internal multi-head attention (hereinafter MHA) mechanisms rather than relying solely on generated text analysis. Operating on already-trained models, the framework employs a two-component architecture: first, a lightweight Long Short-Term Memory (hereinafter LSTM) network processes MHA results in a supervised learning paradigm to perform sub-dataset contribution analysis, enabling precise attribution of generated outputs to specific copyrighted sources such as individual authors’ works; second, a concatenated global projector trained via unsupervised contrastive learning facilitates the detection of non-copyrighted content by learning distinguishing patterns without requiring exhaustive labeled datasets. This methodology addresses critical limitations in existing approaches, which typically apply prompt engineering or semantic classification techniques that treat entire training corpora as uniformly copyrighted and struggle to trace specific source influences.Table below maps the identified challenges to existing solutions that partially address these problems.
@@ 7 Conclusion
To address these limitations, we propose a multilayered filtering pipeline that shifts copyright protection from post-training detection to pre-training prevention. Our framework combines access control mechanisms enhanced with pre-crawl ownership verification, perceptual hashing systems for multimedia content identification, Named Entity Recognition for flagging potential copyrighted sources, machine learning classifiers retrained specifically to distinguish copyrighted from public domain content, and continuous cross-referencing against updated copyright databases. This defense-in-depth approach ensures each layer catches violations missed by others, transforming copyright enforcement from reactive to proactive. By implementing such filtering systems, supported by transparent data provenance documentation and enforced through clear regulatory mandates, can we achieve the dual objectives of protecting creator rights while enabling continued AI innovation.

### v2

@@ 2.4.1 Japan
Japan’s revised Copyright Act (Article 30-4, effective January 1, 2019) allows broad rights to ingest and use copyrighted works for any type of information analysis, including AI training, for both commercial and non-commercial purposes [33]. Unlike the EU, Japan allows commercial use and reportedly even permits training on ”content obtained from illegal sites,” with Japan’s Education Minister indicating AI companies can use ”whatever they want” for training [33].
@@ 2.4.2 Singapore
Singapore’s Copyright Act Amendment of 2021 introduced Section 244, allowing use of copyrighted works for ”computational data analysis” including commercial use, provided users have ”lawful access” to the content [11].
@@ 2.4.3 China
China’s ”Interim Measures for the Management of Generative AI Services” (July 2023) requires that generative AI services ”respect intellectual property rights,” with Article 7.2 stipulating that AI providers must not infringe others’ IP rights [14, 11]. China also requires transparency about training data use and has shown willingness to impose secondary liability on AI platforms for copyright infringement [19, 34].
@@ 2.4.4 South Korea
Korea requires permission for using copyrighted works for AI training, as emphasized in Copyright Commission guidelines published in December 2023. Courts have ruled that web scraping infringes database rights under copyright law [14, 18].
@@ 3.1 Pre-Training License Collection and Content Filtering
Modern Large Language Models (LLMs) require training datasets containing billions of text tokens, images, and other media. The sheer volume of this data collection makes traditional copyright clearance mechanisms practically impossible to implement.
@@
Regulatory Gap: the US Copyright Office noted, ”it is not practically possible to obtain licenses for the volume and diversity of content necessary to power cutting-edge systems” [32]. New websites, blog posts, images, and other copyrighted materials appear online continuously, often before rights holders can implement or update restriction mechanisms. This temporal lag between content publication and rights assertion creates unavoidable exposure windows where AI companies may unknowingly incorporate restricted content into their training datasets.
@@ 3.2 Verification Mechanisms Gaps
Regulatory Gap: On the regulatory side, verification mechanisms remain underdeveloped. The EU AI Code of Practice illustrates this limitation. While Measure 1.4 of the Copyright Chapter [eu2024codeofpractice, Copyright Chapter, Measure 1.4] requires GPAI providers to describe their data sources and data filtering methodologies in technical model documentation, it does not impose any obligation to implement verifiable controls or independent audits of those practices. Nor does it provide technical or procedural standards for assessing whether the documented filtering measures were actually applied or were effective in preventing the use of infringing content.
As a result, compliance is assessed largely on the basis of self-reported documentation. Verification typically occurs only in response to external triggers, such as regulatory complaints or litigation initiated by rights holders, rather than through systematic, proactive oversight. This reflects a predominantly reactive approach to pre-training data governance, in which deficiencies in filtering practices may remain undetected.
@@
An example of mentioned above challenges is The New York Times’ lawsuit against Microsoft and OpenAI [28] which illustrates a serious category of potential copyright infringement. The Times alleges that AI models were trained on wholesale copying of Times articles, and that the resulting AI capabilities directly compete with the newspaper’s core business model. The Times demonstrated instances where AI models could reproduce content substantially similar to specific Times articles, suggesting that training data patterns were being memorized rather than truly transformed. This raises significant problems as the training data used can’t be tracked as well as after training it is impossible to access the dataset.
These three challenges correspond to the open research questions identified in Stanford’s TAIG Database [22] discussed before, demonstrating that the enforcement gaps we have identified reflect broader, unsolved problems in technical AI governance.
@@ 4 Existing Mitigation Strategies
One more solution is the ”pay per crawl” model [3], which allows content creators to have control over who accesses their work. If a creator wants to block all AI crawlers from their content, they should be able to do so. If a creator wants to allow some or all AI crawlers full access to their content for free, they should be able to do that, too. Moreover, pay per crawl grants domain owners full control over their monetization strategy. They can define a flat, per-request price across their entire site, meaning they get paid if an AI company wants to scrape the data from their website. Companies such as CloudFlare [3] have begun implementing and developing this approach, offering platforms for content creators to manage crawler access and monetize their data.
@@
Another solution is Named Entity Recognition (hereinafter NER) [15] which is used in AI Chatbots and security tools. It is a natural language processing technique that identifies and classifies specific entities in text into predefined categories such as: person names, organizations, locations, chemical compounds, works of art. NER could be adapted to identify copyrighted content by: detecting author names, publication titles, or publisher names; identifying trademarked product names or brand references; recognizing specific works (book titles, article headlines, song names); matching against databases of known copyrighted entities. However, NER alone is insufficient for copyright filtering because copyright protects expression, not just entity names. For example, a text could discuss ”The New York Times” without containing any Times content. Conversely, copyrighted text might not contain obvious entity markers, and, moreover, NER can’t detect paraphrased or derivative content.
@@
Also perceptual hashing [24, 25] is a great method that is already in use by many companies. Unlike cryptographic hashes (which change completely with any tiny modification), perceptual hashes create a ”fingerprint” that remains similar even when content is slightly altered (resized, compressed, cropped, color-adjusted). The system can then compare these hashes to identify matching or near-matching content. While this technology is widely deployed by platforms like YouTube and Facebook [31] to identify copyrighted material after upload, it remains underutilized in the AI training data context. AI companies could theoretically check billions of scraped images and videos against copyright databases using perceptual hashing before including them in training datasets, but there’s little evidence of systematic implementation, likely due to computational costs, incomplete database coverage, and legal ambiguity about whether using copyrighted images for training (without reproducing them in outputs) constitutes infringement.
@@
Have I Been Trained [21] is a searchable database that allows artists and creators to check whether their images appear in major publicly documented AI training datasets like LAION-5B and LAION-400M, which were used to train models such as Stable Diffusion. Users can search by uploading images, entering artist names, or providing URLs to discover if their work has been included in training data, and can request opt-out from future dataset versions. However, the tool has significant limitations: it only covers datasets that companies have publicly disclosed (excluding proprietary datasets like OpenAI’s), works retrospectively rather than preventing initial scraping, provides no legal enforcement mechanism, covers only images rather than text or code, and relies entirely on voluntary compliance from AI companies. This shows the reactive nature of current copyright protection mechanisms, creators can discover violations only after their work has already been scraped and used for training, and opt-out requests don’t affect models already trained on the data, highlighting the gap between post-collection transparency tools and the proactive pre-training filtering mechanisms needed to prevent copyright infringement before it occurs.
@@
InnerProbe [17] presents a post-training framework for identifying copyright-related influences in Large Language Model outputs by leveraging internal multi-head attention (hereinafter MHA) mechanisms rather than relying solely on generated text analysis. Operating on already-trained models, the framework employs a two-component architecture: first, a lightweight Long Short-Term Memory (hereinafter LSTM) network processes MHA results in a supervised learning paradigm to perform sub-dataset contribution analysis, enabling precise attribution of generated outputs to specific copyrighted sources such as individual authors’ works; second, a concatenated global projector trained via unsupervised contrastive learning facilitates the detection of non-copyrighted content by learning distinguishing patterns without requiring exhaustive labeled datasets. This methodology addresses critical limitations in existing approaches, which typically apply prompt engineering or semantic classification techniques that treat entire training corpora as uniformly copyrighted and struggle to trace specific source influences.Table below maps the identified challenges to existing solutions that partially address these problems.
@@ 7 Conclusion
To address these limitations, we propose a multilayered filtering pipeline that shifts copyright protection from post-training detection to pre-training prevention. Our framework combines access control mechanisms enhanced with pre-crawl ownership verification, perceptual hashing systems for multimedia content identification, Named Entity Recognition for flagging potential copyrighted sources, machine learning classifiers retrained specifically to distinguish copyrighted from public domain content, and continuous cross-referencing against updated copyright databases. This defense-in-depth approach ensures each layer catches violations missed by others, transforming copyright enforcement from reactive to proactive. By implementing such filtering systems, supported by transparent data provenance documentation and enforced through clear regulatory mandates, we can achieve the dual objectives of protecting creator rights while enabling continued AI innovation.
Acknowledgement
The authors would like to thank Hannah Khier for his time and effort in reviewing this manuscript and for providing valuable comments and suggestions that helped improve its clarity and quality.